OpenStudy (anonymous):

MEDAL! What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8 A. neither parallel nor perpendicular B. parallel C. perpendicular D. they are the same line

OpenStudy (anonymous):

slope of first line is m1=-1/2 slope of second line is m2=2 since m1*m2=-1 so they are perpendicular .

OpenStudy (anonymous):

Thankyou!!! :D